OpenStudy (anonymous):

x=sin^2t and y=sin^4t to rectangular

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aum):

\[ x = \sin^2(t) \\ y = \sin^4(t) \\ x^2 = \sin^4(t) = y\\ y = x^2 \]

OpenStudy (anonymous):

aum please help i beg u http://openstudy.com/updates/5473e74ee4b0db93229f17a5

OpenStudy (anonymous):

help :D

OpenStudy (anonymous):

@satellite73